OpenStudy (babyslapmafro):

Please help, I need to know how to solve a problem like this for an exam tomorrow: Find the dimensions of the rectangle with the largest area that can be inscribed under the curve y=cos(x). (Click to see parameters and drawing)

OpenStudy (babyslapmafro):

\[0 \le x \le \frac{ \pi }{ 2 }\]

OpenStudy (babyslapmafro):



OpenStudy (anonymous):

Is this for calculus (derivatives) or something else?

OpenStudy (babyslapmafro):

Sorry, I should have made that clear. We are covering absolute and relative max/min.

OpenStudy (anonymous):

So this is Alg. II?

OpenStudy (babyslapmafro):

Calc 1

OpenStudy (anonymous):

Ok, so assuming that you can take a derivative, you can create a function describing the area of the rectangle formed, and then take the derivative, set that function to zero and find the relative max. that way. Note that given a function with a relative max/min, that max/min can be found by taking the derivative of that function and setting that derivative to zero and solving.

OpenStudy (babyslapmafro):

Ok, how to I create a function that describes the area of the rectangle?

OpenStudy (anonymous):

What is the height of the rectangle as a function of what they give you (the angle or the cosine of the angle) and then what is the width of the rectangle as a function of what they give you (the angle or the cosine of the angle?)